OpenStudy (anonymous):

I need help with radical expressions. the question is radical (a+5)^2

OpenStudy (anonymous):

\[\sqrt{b^2} = |b|\]

OpenStudy (anonymous):

In this case your b is a+5, but the property still holds.

OpenStudy (anonymous):

so how do i solve the problem. I am confused it says you use any real number for the variable. So if i would use 2 would i take 5*5 and then add 2 to get the answer?

OpenStudy (anonymous):

a can be any real number. You don't just pick one.

OpenStudy (anonymous):

Just re-write the expression as I did, but instead of b use a+5.

OpenStudy (anonymous):

okay....i have more would you be able to help me with them? i do not understand any radicals.

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

sounds good here they are! \[(\sqrt{7})(-2\sqrt{7+4})\] \[\sqrt[5]{x}^{10}\] \[\sqrt[3]{-27x ^{3}}\]

OpenStudy (anonymous):

the bottom two are easier. Remember this rule (write it down): \[\sqrt[r]{a^bc^k} = a^{b/r}c^{k/r}\]

OpenStudy (anonymous):

you should be able to use it immediately to find the second problem

OpenStudy (anonymous):

to find the answer to the 3rd, re-write -27 as a power of 3.

OpenStudy (anonymous):

err a power of -3 rather

OpenStudy (anonymous):

okay great! thank you

OpenStudy (anonymous):

once you are done with them we can take a stab at the first one.

OpenStudy (anonymous):

okay i did them

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I do not think i did them correctly but just really want to be done so i just put x^2 and \[\sqrt[3]{3^{3}}\]

OpenStudy (anonymous):

x^2 was correct. The second is not though. \[\sqrt[3]{-27x^3} = \sqrt[3]{(-3)^3x^3} = (-3)^{3/3}x^{3/3} = (-3)^1x^1 = -3x\]

OpenStudy (anonymous):

hm..see how comes it seems so easy for you lol. i really hate algebra

OpenStudy (anonymous):

it just takes practice. If you got stuck you should have just asked for clarification ;)

OpenStudy (anonymous):

do i have to put the answer with the radical sign?

OpenStudy (anonymous):

what radical sign? Is there a radical in the answer?

OpenStudy (anonymous):

well some times clarification does not always make me understand it any better lol :)

OpenStudy (anonymous):

\[\sqrt{?}\] isnt that a radical sign?

OpenStudy (anonymous):

yes it is, do you see one in my answer?

OpenStudy (anonymous):

no but didnt know if I had to put it in or not? guess not